Practice Test Question.
Log 3 (x-2)= -1

Answer by stanbon(75887)

Log 3 (x-2)= -1
---
x-2 = 3^-1
---
x-2 = 1/3
---
x = 2 1/3
==================
cheers,
Stan H.